# CriteoPrivateAd: A Real-World Bidding Dataset to Design Private Advertising Systems

### v2

@@ Abstract
In the past years, many proposals have emerged in order to address online advertising use-cases without access to third-party cookies. All these proposals leverage some privacy-enhancing technologies such as aggregation or differential privacy. Yet, no public and rich-enough ground truth is currently available to assess the relevancy of aforementioned private advertising frameworks. We are releasing the largest, in terms of number of features, bidding dataset specifically built in alignment with the design of major browser vendors proposals such as Chrome Privacy Sandbox. This dataset, coined CriteoPrivateAd, stands for an anonymised version of Criteo production logs and provides sufficient data to learn bidding models commonly used in online advertising under many privacy constraints (delayed reports, display and user-level differential privacy, user signal quantisation or aggregated reports). We ensured that this dataset, while being anonymised, is able to provide offline results close to production performance of adtech companies including Criteo - making it a relevant ground truth to design private advertising systems. The dataset is available in Hugging Face: https://huggingface.co/datasets/criteo/CriteoPrivateAd.
@@ 1 Introduction
Contributions. Albeit many private advertising proposals have emerged, the advertising industry is currently lacking a common ground truth in order to assess the privacy/utility trade-off of such envisioned private advertising systems. A public dataset, using synthetic or anonymised industrial data, would allow benchmarking of privacy-preserving mechanisms, fostering transparency, innovation, and ensuring effective solutions for user data protection and online advertising use-cases. This paper is an attempt to fill this gap. We are open-sourcing CriteoPrivateAd: the largest real-world anonymised bidding dataset in terms of number of features from Criteo production data. Leveraging this dataset, researchers and engineers would be able to:
@@ 2 Related Work
CriteoPrivateAd is versatile enough to be leveraged to assess the offline performance of both non-private and private advertising approaches. Regarding the latter, benchmarks of bidding models based on differential privacy (DP), or other obfuscation techniques such as aggregation and signal quantisation (e.g. restricting user signals to a few bits) are of relevant interest, being both studied in the academic and industrial research and engineering communities. Since DP stands for the workhorse methodology to derive private training frameworks, we will focus primarily on DP and show how this open-source release is necessary to benchmark private advertising systems with respect to current open-source bidding datasets.
@@ 3 Background on Differential Privacy
4 CriteoPrivateAd Dataset
This section describes how the bidding dataset CriteoPrivateAd has been built, including business risk mitigations that were taken to preserve Criteo intellectual property; and how it can be leveraged to assess offline performance close to that of real-world production advertising systems. The dataset is available in Hugging Face: https://huggingface.co/datasets/criteo/CriteoPrivateAd.
@@ 4.1 Description
Overview. The CriteoPrivateAd dataset contains 100 millions displays spanning 30 consecutive days of data and more than 100 relevant features to learn common bidding models. Indeed, we are providing binary labels such as click or landed click; as well as integer labels such as the number of sales attributed to a specific display. Such labels are retrieved from third-party cookie traffic on Chrome and hence not necessarily representative of Chrome Protected Audiences traffic. These choices have been made in order to have a sufficient number of positive examples enabling to learn relevant bidding models with a performance close to that of production systems. Beyond features and labels relevant for training bidding models, we are also providing information required to simulate Privacy Sandbox APIs and emulate user-level DP, namely: hashed user ids, hashed publisher ids, hashed campaign ids (could be used as a proxy of advertiser ids) and time delta between the bid request timestamp and the click or sale event. This dataset is partionned by day to facilitate exploration, model seeding and train/validation/test split.
@@
Reproducing Production Offline Performance. We decided to sub-sample negative examples to ensure that the number of clicks remains sufficient for achieving representative performances, leading to a click-through rate CTR $\approx 0.36$ and a conversion rate CVR $\approx 0.02$, where a conversion here means a positive sale event (i.e., at least one product has been bought by the user). Hence, it is important to rescale validation metrics for interpretable results, close to production performance. Let $p_{0}$ be the CTR in CriteoPrivateAd, $p^{\star}$ the targeted CTR representative of online performance. By denoting $y\in\{0,1\}$ be the true label $y$, $\hat{y}\in[0,1]$ the prediction and $\ell$ the loss, we must use for validation the re-weighted loss (Elkan, 2001, Theorem 1):
@@ 4.2 Column Semantics
  • uuid: user id consistent over the day. The same user will have two different user ids for two different days. To clarify the difference between uuid and a device id, note that the data in CriteoPrivateAd comes mainly from the third-party cookie in the Chrome instance, and the rest from aggregated data from the 3rd-party cookie and app data from the same device (cookie id is matched with gaid from apps).
@@ 4.3 Baseline experiments
The CriteoPrivateAd dataset provides sufficient number of features and displays to learn an algorithm with relevant offline performance, ensuring its relevance for industrial benchmarks and applications. In Table 2, we provide results obtained on this dataset using baseline models based on logistic regression, with a given targeted CTR and the associated re-weighted loss, as defined in (1).
@@ 5 Private Bidding Optimisation Paradigms
To ease the use of CriteoPrivateAd, we are providing in this section some brief private training paradigms to learn bidding models, that are currently considered or researched in the advertising industry. As before, these use-cases are mainly inspired from the Chrome Privacy Sandbox but some of them are also aligned with where the industry is leaning. Note that these use-cases are obviously not exhaustive and that CriteoPrivateAd is versatile enough to be used in many applications.
@@ 5.1 User vs Display-level DP
This dataset being sampled at the display level, it can be naturally used to design and test private bidding model training strategies based on display-level differential privacy (also coined item-level DP in the academic literature). Albeit we are providing a user identifier, this dataset cannot be used directly, without involving some bias, to assess the relevancy of user-level DP frameworks. Indeed, since we sampled on a display basis and not on a user one, the distribution of user’s contributions in CriteoPrivateAd is different from the online traffic one. To tackle this issue, we are providing two ways to use this dataset in a user-level DP context.
CriteoPrivateAd might be representative to what is sent to a private aggregation server. In the Privacy Sandbox context, the report collector (e.g., the adtech company willing to collect event-level reports to train bidding models) will have access to display-level reports (either encrypted or not), sampling per display could be performed by the report collector, which will then send the sampled batch of reports to a private aggregation server (e.g., a Private Aggregation service). This server will then implement a user-level DP framework, by for instance capping the number of displays per user (with some fixed or flexible logic, e.g., by sampling uniformly all user’s displays). The final dataset will then be used for aggregation or training purposes. To avoid any bias issue between offline training and online inference, we will need the sampling ratio of negatives and positives to align the training loss with the real data distribution.
Figure 1: Discrepancy between empirical distributions of users’ contributions in CriteoPrivateAd and associated online traffic.
CriteoPrivateAd could be aligned to the true user distribution. In addition to CriteoPrivateAd dataset, we will provide the data used to plot the two histograms depicted in Figure 1. To simulate what happens on the original non-sampled dataset, we propose a method to build a semi-synthetic dataset from CriteoPrivateAd, where the distribution of displays per user matches the original non-sampled dataset. By denoting $p$ the real distribution (depicted in blue in the above figure) and $q$ the approximate one (depicted in orange), we can build importance ratios $p/q$ per bucket and use them to sub-sample CriteoPrivateAd data to match $p$. Naive sampling might lead to duplicated (if we fix a too large dataset size) or very few rows (if we sub-sampled directly each bucket) in the final dataset. To avoid this issue, we can take all displays associated to the last bucket in CriteoPrivateAd (e.g., number of displays per user = 20 if the support is [1,20]), use $p$ to determine a dataset size (i.e., dataset size = nb_users_having_20_displays divided by Proba(nb_displays_per_user = 20); and sample CriteoPrivateAd based on $p$.
@@ 5.3 Aggregation API - Learning from Label Proportions
Assuming that a feature vector $x$ lies in $\mathcal{X}$, and that labels are binary, i.e., $y\in\{0,1\}$, given $\mathbf{A}\in\mathcal{X}$ the Private Aggregation Service of Chrome Privacy Sandbox allows to get a noisy estimate of $\mathbb{P}\left(y=1|\mathbf{A}\right)$, by defining the key $k(x,\mathbf{A})=1$ if and only if $x\in\mathbf{A}$. In other words, by using this Aggregation Service, we get noisy label proportions, where the noise is applied on the label proportion $\mathbb{P}\left(y=1|\mathbf{A}\right)$ using global differential privacy. As of today, the Laplace mechanism is used in Chrome Privacy Sandbox, see Definition 2. The set $\mathbf{A}$ stands for the so-called ”bag” in the learning from label proportions paradigm (Quadrianto et al., 2008). This paradigm can be simulated using CriteoPrivateAd by aggregating display-level data and noising the associated label proportions.
@@ 6 Conclusion
We release in an open-source manner CriteoPrivateAd, the largest and richest bidding dataset to learn a variety of click and conversion models. This dataset can be used to learn standard bidding models but its structure has been designed to fit browser vendors private advertising proposals and associated constraints. Through this open-source release, we hope to foster innovation in both the academic and industrial communities to design scalable private bidding optimisation strategies which help protecting users’ data while ensuring an economically viable open internet.

### v3

# CriteoPrivateAds: A Real-World Bidding Dataset to Design Private Advertising Systems

@@ Abstract
In the past years, many proposals have emerged in order to address online advertising use-cases without access to third-party cookies. All these proposals leverage some privacy-enhancing technologies such as aggregation or differential privacy. Yet, no public and rich-enough ground truth is currently available to assess the relevancy of aforementioned private advertising frameworks. We are releasing the largest, in terms of number of features, bidding dataset specifically built in alignment with the design of major browser vendors proposals such as Chrome Privacy Sandbox. This dataset, coined CriteoPrivateAds, stands for an anonymised version of Criteo production logs and provides sufficient data to learn bidding models commonly used in online advertising under many privacy constraints (delayed reports, display and user-level differential privacy, user signal quantisation or aggregated reports). We ensured that this dataset, while being anonymised, is able to provide offline results close to production performance of adtech companies including Criteo - making it a relevant ground truth to design private advertising systems. The dataset is available in Hugging Face: https://huggingface.co/datasets/criteo/CriteoPrivateAd.
@@ 1 Introduction
Contributions. Albeit many private advertising proposals have emerged, the advertising industry is currently lacking a common ground truth in order to assess the privacy/utility trade-off of such envisioned private advertising systems. A public dataset, using synthetic or anonymised industrial data, would allow benchmarking of privacy-preserving mechanisms, fostering transparency, innovation, and ensuring effective solutions for user data protection and online advertising use-cases. This paper is an attempt to fill this gap. We are open-sourcing CriteoPrivateAds: the largest real-world anonymised bidding dataset in terms of number of features from Criteo production data. Leveraging this dataset, researchers and engineers would be able to:
@@ 2 Related Work
CriteoPrivateAds is versatile enough to be leveraged to assess the offline performance of both non-private and private advertising approaches. Regarding the latter, benchmarks of bidding models based on differential privacy (DP), or other obfuscation techniques such as aggregation and signal quantisation (e.g. restricting user signals to a few bits) are of relevant interest, being both studied in the academic and industrial research and engineering communities. Since DP stands for the workhorse methodology to derive private training frameworks, we will focus primarily on DP and show how this open-source release is necessary to benchmark private advertising systems with respect to current open-source bidding datasets.
@@ 3 Background on Differential Privacy
4 CriteoPrivateAds Dataset
This section describes how the bidding dataset CriteoPrivateAds has been built, including business risk mitigations that were taken to preserve Criteo intellectual property; and how it can be leveraged to assess offline performance close to that of real-world production advertising systems. The dataset is available in Hugging Face: https://huggingface.co/datasets/criteo/CriteoPrivateAd.
@@ 4.1 Description
Overview. The CriteoPrivateAds dataset contains 100 millions displays spanning 30 consecutive days of data and more than 100 relevant features to learn common bidding models. Indeed, we are providing binary labels such as click or landed click; as well as integer labels such as the number of sales attributed to a specific display. Such labels are retrieved from third-party cookie traffic on Chrome and hence not necessarily representative of Chrome Protected Audiences traffic. These choices have been made in order to have a sufficient number of positive examples enabling to learn relevant bidding models with a performance close to that of production systems. Beyond features and labels relevant for training bidding models, we are also providing information required to simulate Privacy Sandbox APIs and emulate user-level DP, namely: hashed user ids, hashed publisher ids, hashed campaign ids (could be used as a proxy of advertiser ids) and time delta between the bid request timestamp and the click or sale event. This dataset is partionned by day to facilitate exploration, model seeding and train/validation/test split.
@@
Reproducing Production Offline Performance. We decided to sub-sample negative examples to ensure that the number of clicks remains sufficient for achieving representative performances, leading to a click-through rate CTR $\approx 0.36$ and a conversion rate CVR $\approx 0.02$, where a conversion here means a positive sale event (i.e., at least one product has been bought by the user). Hence, it is important to rescale validation metrics for interpretable results, close to production performance. Let $p_{0}$ be the CTR in CriteoPrivateAds, $p^{\star}$ the targeted CTR representative of online performance. By denoting $y\in\{0,1\}$ be the true label $y$, $\hat{y}\in[0,1]$ the prediction and $\ell$ the loss, we must use for validation the re-weighted loss (Elkan, 2001, Theorem 1):
@@ 4.2 Column Semantics
  • uuid: user id consistent over the day. The same user will have two different user ids for two different days. To clarify the difference between uuid and a device id, note that the data in CriteoPrivateAds comes mainly from the third-party cookie in the Chrome instance, and the rest from aggregated data from the 3rd-party cookie and app data from the same device (cookie id is matched with gaid from apps).
@@ 4.3 Baseline experiments
The CriteoPrivateAds dataset provides sufficient number of features and displays to learn an algorithm with relevant offline performance, ensuring its relevance for industrial benchmarks and applications. In Table 2, we provide results obtained on this dataset using baseline models based on logistic regression, with a given targeted CTR and the associated re-weighted loss, as defined in (1).
@@ 5 Private Bidding Optimisation Paradigms
To ease the use of CriteoPrivateAds, we are providing in this section some brief private training paradigms to learn bidding models, that are currently considered or researched in the advertising industry. As before, these use-cases are mainly inspired from the Chrome Privacy Sandbox but some of them are also aligned with where the industry is leaning. Note that these use-cases are obviously not exhaustive and that CriteoPrivateAds is versatile enough to be used in many applications.
@@ 5.1 User vs Display-level DP
This dataset being sampled at the display level, it can be naturally used to design and test private bidding model training strategies based on display-level differential privacy (also coined item-level DP in the academic literature). Albeit we are providing a user identifier, this dataset cannot be used directly, without involving some bias, to assess the relevancy of user-level DP frameworks. Indeed, since we sampled on a display basis and not on a user one, the distribution of user’s contributions in CriteoPrivateAds is different from the online traffic one. To tackle this issue, we are providing two ways to use this dataset in a user-level DP context.
CriteoPrivateAds might be representative to what is sent to a private aggregation server. In the Privacy Sandbox context, the report collector (e.g., the adtech company willing to collect event-level reports to train bidding models) will have access to display-level reports (either encrypted or not), sampling per display could be performed by the report collector, which will then send the sampled batch of reports to a private aggregation server (e.g., a Private Aggregation service). This server will then implement a user-level DP framework, by for instance capping the number of displays per user (with some fixed or flexible logic, e.g., by sampling uniformly all user’s displays). The final dataset will then be used for aggregation or training purposes. To avoid any bias issue between offline training and online inference, we will need the sampling ratio of negatives and positives to align the training loss with the real data distribution.
Figure 1: Discrepancy between empirical distributions of users’ contributions in CriteoPrivateAds and associated online traffic.
CriteoPrivateAds could be aligned to the true user distribution. In addition to CriteoPrivateAds dataset, we will provide the data used to plot the two histograms depicted in Figure 1. To simulate what happens on the original non-sampled dataset, we propose a method to build a semi-synthetic dataset from CriteoPrivateAds, where the distribution of displays per user matches the original non-sampled dataset. By denoting $p$ the real distribution (depicted in blue in the above figure) and $q$ the approximate one (depicted in orange), we can build importance ratios $p/q$ per bucket and use them to sub-sample CriteoPrivateAds data to match $p$. Naive sampling might lead to duplicated (if we fix a too large dataset size) or very few rows (if we sub-sampled directly each bucket) in the final dataset. To avoid this issue, we can take all displays associated to the last bucket in CriteoPrivateAds (e.g., number of displays per user = 20 if the support is [1,20]), use $p$ to determine a dataset size (i.e., dataset size = nb_users_having_20_displays divided by Proba(nb_displays_per_user = 20); and sample CriteoPrivateAds based on $p$.
@@ 5.3 Aggregation API - Learning from Label Proportions
Assuming that a feature vector $x$ lies in $\mathcal{X}$, and that labels are binary, i.e., $y\in\{0,1\}$, given $\mathbf{A}\in\mathcal{X}$ the Private Aggregation Service of Chrome Privacy Sandbox allows to get a noisy estimate of $\mathbb{P}\left(y=1|\mathbf{A}\right)$, by defining the key $k(x,\mathbf{A})=1$ if and only if $x\in\mathbf{A}$. In other words, by using this Aggregation Service, we get noisy label proportions, where the noise is applied on the label proportion $\mathbb{P}\left(y=1|\mathbf{A}\right)$ using global differential privacy. As of today, the Laplace mechanism is used in Chrome Privacy Sandbox, see Definition 2. The set $\mathbf{A}$ stands for the so-called ”bag” in the learning from label proportions paradigm (Quadrianto et al., 2008). This paradigm can be simulated using CriteoPrivateAds by aggregating display-level data and noising the associated label proportions.
@@ 6 Conclusion
We release in an open-source manner CriteoPrivateAds, the largest and richest bidding dataset to learn a variety of click and conversion models. This dataset can be used to learn standard bidding models but its structure has been designed to fit browser vendors private advertising proposals and associated constraints. Through this open-source release, we hope to foster innovation in both the academic and industrial communities to design scalable private bidding optimisation strategies which help protecting users’ data while ensuring an economically viable open internet.